I have to find out which bag of food is the better buy. I tried to solve the problem, but I don't think it is correct Can you check?
Weight Bag Size (radius x height) Cost per Bag
4 lb bag 4.5x11 $10.88
15 lb bag 6.5x15 $29.69
35 lb bag 9x23 $52.46

Answer by richwmiller(17219)

Size of the bag has nothing to do with cost efficiency.
Price per pound is what you want.
10.88/4.5=2.4177 $2.42 per pound
29.69/15=1.979=$1.98 per lb.
52.46/35=1.4988=$1.50 per lb.